OpenStudy (anonymous):

Let \(\{a_n\}\mid\sum\sqrt{a_n}<\infty\) and \(\{b_n\}\in\mathbb R\). Now:\[E=\left\{\frac{a_n}{\left|x-b_n\right|}<\infty\right\}\]Prove that \(\mathbb R-E=\emptyset\).

OpenStudy (anonymous):

@Zarkon @FoolForMath @KingGeorge @TuringTest :3 don't mind if I drag you in.

OpenStudy (kinggeorge):

Just to be clear, \(\emptyset\) is denoting the empty set correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Made a correction. :\ Minor detail, but important.

OpenStudy (kinggeorge):

I suppose we also have to consider complex sequences of \(\{a_n\}\)?

OpenStudy (anonymous):

I think so.

OpenStudy (kinggeorge):

In that case, I have no idea whatsoever.

OpenStudy (anonymous):

If we were to restrict \(\{a_n\}\in\mathbb R\), what would you do? I still don't know what I'd do about this.

OpenStudy (kinggeorge):

Even then, I'm still not sure where to go. If we were to restrict everything to the positive reals, I might be able to get somewhere, but I'm still not sure where to start.

OpenStudy (anonymous):

Well, thanks for the attempt. :)

OpenStudy (anonymous):

Are the sequences \[ a_n\] and \[b_n\] fixed and we consider the set E of x in R so the the inequality is satisfied?

OpenStudy (anonymous):

@eliassaab Yes.

OpenStudy (anonymous):

Take \[a_n =\frac 1 {n^4}\\ b_n =0,\text { for all }n \] The number x=0 does not belong to E. So R-E is not the empty set.

OpenStudy (anonymous):

Alright, cheers.